OpenStudy (anonymous):

Please help !! How would you find the 8th term of the arithmetic sequence: -3.2, -3.4, -3.6, -3.8, ...

OpenStudy (ivettef365):

use formula an+b a = sequence n = nth term b = difference from 1st# and sequence

OpenStudy (ivettef365):

I show you how to do this a = -0.2 n = 8 b = -3.2 - (-0.2) = -3 now plug on formula

OpenStudy (ivettef365):

do you understand what I did ?

OpenStudy (anonymous):

how did you get the values for a & n ?

OpenStudy (anonymous):

i have never seen it done that way until now wow

OpenStudy (ivettef365):

the sequence is adding -0.2 to each number -3.2 + -0.2 = -3.4 +-0.2 = -3.6 and so on

OpenStudy (ivettef365):

the question ask for the 8th term, that is what you are looking for so n = 8

OpenStudy (anonymous):

okay i see. so every time the a value would be -0.2 ?

OpenStudy (ivettef365):

so if you plug the numbers this is what you get: an + b -0.2 (8) + (-3) -1.60 -3 = -4.6

OpenStudy (anonymous):

how would you have done it ? @nocontact115

OpenStudy (ivettef365):

no every time, just on this specific problem

OpenStudy (anonymous):

say if there were a different problem how would you find it ?

OpenStudy (ivettef365):

you can add -0.2 until you get to the 8th term

OpenStudy (anonymous):

got it ... thanks(:

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

@ idontgetit i would have done it by using subtraction to find the difference between the numbers and added it to keep the pattern going.